Circular No. 7950

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2002 P1 S. Pravdo, K. Lawrence and E. Helin, Jet Propulsion Laboratory, re-

port the NEAT team’s discovery with the 1.2-m Schmidt telescope at Palomar of a comet with a 7′′ tail in p.a. 235^{o}. Confirmation of the object’s diffuse appearance has been provided by G. Hug, Farpoint Observatory:

2002 UT α_{2000} δ_{2000} m_{1} Observer
Aug. 7.38888 21^{h}16^{m}06.^{s}93 −9^{o}47′55′′.3 19.6 NEAT
7.40417 21 16 06.50 −9 47 54.2 19.2  _{”}
7.41704 21 16 06.16 −9 47 52.8 19.9  _{”}
8.25898 21 15 44.62 −9 46 50.0 18.3 Hug
8.27350 21 15 44.37 −9 46 47.7 18.4  _{”}

V4334 SAGITTARII R. W. Russell, D. K. Lynch and D. L. Kim, The Aerospace Corpo-

ration; M. L. Sitko, University of Cincinnati; E. Polomski, University of Minnesota; and H. B. Hammel, Space Science Institute, report 3–14-μm spectroscopy of V4334 Sgr (Sakurai’s object) using BASS on the NASA IRTF on July 31.38 UT: “The spectrum is well fitted by a 430(±10)-K blackbody. The data appear consistent with a very weak (possibly silicate) absorption feature near 10–11 μm. Monochromatic magnitudes [3.5 μm], [4.8 μm] and [10.2 μm] were 5.09, 2.64 and 0.45, respectively, all ±0.05 or less. Observations taken with The Aerospace Corporation’s near-infrared (0.8–2.5 μm) imaging spectrograph at the Lick Observatory on July 18.3 by R. J. Rudy, D. K. Lynch, S. M. Mazuk, C. C. Venturini, R. C. Puetter and R. B. Perry reveal a featureless continuum rising rapidly at longer wavelengths and joining smoothly with the BASS/IRTF spectrum. H and K short magnitudes were 15.8 and 11.2, respectively, both ±0.2. Compared to the 1998 observations (Lynch et al. 2002, Astrophys. Space Sci. 279, 57), the object’s near-infrared flux has dropped dramatically, but the infrared luminosity has increased by ∼ 40 percent.”

COMET C/2002 O4 (HÖNIG) Sitko, Lynch, Russell, Hammel and Polomski report similar spectros-

copy of comet C/2002 O4 on Aug. 1.54 UT: “The infrared flux peaked near 10 μm, suggesting a temperature of ∼ 280 ± 20 K (equilibrium blackbody temperature = 243 K). There was a silicate-emission feature between ∼ 8.5 and 11.5 μm extending ∼ 20 percent above the continuum. Narrow-band [M] and [N] magnitudes were 9.7 and 4.4, respectively, both ±0.1.”

2002 August 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden